A wire is Bent into a circle of radius a and a total charge Q is distributed uniformly around it.find the magnetic field at the center if the wire rotates with an angular velocity w about the axis of the circle

1
Expert's answer
2022-01-24T11:17:38-0500

Given charge q, rotating in a circular motion of radius a


Linear charge density, "\lambda=\frac{q}{2 \pi a}"


Time period, "T= \frac{2 \pi}{w}"



Current in circular motion, "I=\frac{q}{T}=\frac{qw}{2 \pi}"


Magntic field, "B= \frac{\mu_oqw}{2a}=\frac{\mu_oqw}{2a}×\frac{ q}{2 \pi a}=\frac{\mu_oqw}{4\pi a}"


Magnetic field at the centre is "\frac{\mu_oqw}{4\pi a}"